A person in a kayak starts paddling, and it accelerates from 0 to 0.515 m/s in a distance of 0.319 m. If the combined mass of the person and the kayak is 64.2 kg, what is the magnitude of the net force acting on the kayak?
Answer #1

using the equation of linear motion we have

v^2 - u^2= 2aS

0.515^2 - 0 = 2*a*0.319

a = 0.4157 m/s^2

force = mass*acceleration

= ma

= 64.2 * 0.4157 = 26.68 N